Scipio Garage has a debt to asset ratio of 40 percent, sales of $8,000, net income of $600, and total debt of $2,400. What is its return on equity. Enter you answer in whole numbers. For example, enter 10% as 10.
Answer #1

Debt/asset = 0.4

Asset = 2400/0.4 = 6000

Equity = asset-debt = 6000-2400 = 3600

ROE = net income/equity = 600/3600 = 16.67%